OpenStudy (anonymous):

Ball is fired at 112ft/sec. H= -16t2^+112t+30. What is maximum height and time to achieve that height. Cannot calculate time?

OpenStudy (atlas):

If you state your complete question - maybe I can help you

OpenStudy (tkhunny):

?? Have you considered t = -112/(2*(-16)) = 56/16 = 28/8 = 14/4 = 7/2?

OpenStudy (gebooors):

Derive H(t) and solve when H'(t) = 0. Then you get time when ball is in the highest point.

OpenStudy (anonymous):

solving for h'(t) = 0 is the exact same thing as just using -b/2a

OpenStudy (anonymous):

Thanks, done.